# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

### v2

@@ Abstract
Networks of physical units can vary from a stationary set of spatially-embedded links to a collection of mobile agents that undergo transient social interactions. In living cells, mitochondria form architectures that span across these regimes, transitioning between fragmented, partly connected, and highly fused structures depending on cell type and state. Diffusive transport of biomolecular components through these networks limits the heterogeneity of the mitochondrial population. Here we address the connection between dynamic network architecture and the rate of diffusive mixing through simulations and analytic models that incorporate fusion, fission, and rearrangement. We find that the material delivered from a source to the rest of the network depends on the network dimensionality and a balance of competing timescales for encounter, fusion, and diffusive dispersion. We extract morphological and dynamic parameters for mitochondrial networks in three human cell lines, demonstrating that different cells span across both the physical and social network regimes. Mixing in mitochondrial networks is shown to be limited by material transport through connected tubules for slowly-diffusing particles and by inter-mitochondrial encounter rates for rapidly diffusing ones. These results provide a quantitative basis for predicting the homogenization of proteins, lipids, ions, or genetic material through the mitochondrial population. The general principles identified in this work capture diffusive spreading through both social and physical networks, unifying a continuum of spatial network architectures.
Introduction
@@
Diffusive transport through networks has been studied in a variety of contexts, including disease spread in epidemiology [1, 2], innovations in social networks [3], communication among insects [4, 5], diffusion of signals and nutrients in the brain extracellular space [6, 7], and oil recovery in porous rock [8]. These systems can be separated into two broad classes: one where the topology of network connections is stationary and limited to nearest neighbors, and another where network nodes are mobile and promiscuous, interacting with many different partners over time.
@@
Physical networks form an important category of stationary, spatially-constrained network structures [9]. In such networks, edges represent objects subject to physical limitations (space-filling, steric repulsion, etc.) [10, 11] connecting degree-1 tips and degree-3 junctions, with higher degree nodes exceedingly rare [12, 13]. Examples include fungal mycelia [14], neuronal synaptic networks [15], and interconnected pores in rocks [16]. Transport behaviors on stationary networks can be described by the graph Laplacian, whose eigenvalues govern the spreading timescales [17, 18, 19]. In some percolation problems, physical networks become dynamic as edges are allowed to flicker between an active and inactive state [20, 21].
@@
A distinct set of approaches considers signal transmission in ‘social’ networks of transiently-interacting mobile units. In such networks, the mixing behavior is defined by distributions of contact durations and inter-contact times [4, 22], which may be constrained by spatial embedding [2, 1]. Notably, many previously studied social network systems focus on the spreading of a non-diluting signal (infection, information, etc.) [4, 2, 3], in contrast to dispersion of mass-conserving physical material [23, 21, 24].
@@
Despite the distinct modeling approaches employed, physical and social networks lie on a continuum of temporal network structures [22] with varying timescales of topological rearrangement. The intracellular environment exemplifies this spectrum. The endoplasmic reticulum forms a well-connected physical lattice of membrane-bound tubules that enables diffusive transport of proteins [24] and ions [25] across the entire cell. At the opposite extreme, the population of endocytic vesicles constitutes a social network of discrete transiently-interacting compartments [26, 27].
Mitochondria form another intracellular network with striking structural variability. Mitochondrial network architectures vary across cell types [28, 29, 30, 31], and transition between fragmented and hyperfused states in response to disease [32, 33, 34, 35, 36, 37], metabolic conditions [38, 39, 40], calcium signalling and apoptosis [41, 42], and cell cycle progression [43, 44]. These structural changes are thought to support important functions such as genetic complementation [45, 46], dilution of harmful reactive oxygen species (ROS) [47], power cabling [48, 49], quality control [50, 51, 52, 53], and modulation of mitochondrial heterogeneity across the population [54, 55]. These cellular functions rely on the connectivity-dependent diffusive spread of proteins, lipids, and/or ions through the underlying mitochondrial network [56, 21].
@@
To elucidate how mitochondrial structure and dynamics govern the mixing of effectively diffusive particles, we introduce a quantitative framework for transport on spatially-embedded dynamic networks. We show that network connectivity governs spreading behavior by modulating several competing timescales: the cluster filling rate, encounter rate, and fusion rate, as well as the decay rate of the spreading material. In particular, we demonstrate a transition from rapid but low-dimensional spread in well-connected networks to slower three-dimensional spreading through an interacting population. Due to the different dimensionalities involved, particles with distinct diffusivities and networks with distinct structures can exhibit qualitatively different scaling behavior for dispersion from a source. We use the model to predict spreading rates for mitochondrial networks in three different human cell types, demonstrating the structural diversity and different regimes of relevance for these networks.
@@
Figure 1: Dynamic network simulation framework. (a) A network is formed by interacting units which diffuse through 3D space with diffusivity $D_{1}$. Nearby units can undergo tip-tip and tip-side fusion with rates $k_{u1},k_{u2}$ while connected nodes undergo fission at rate $k_{f},\frac{3}{2}k_{f}$, respectively. Material spreads along connected units with diffusivity $D_{p}$ and decays over time with rate $k_{d}$. Red indicates the concentration of material in a given unit. (b) Snapshots of spreading on simulated networks at steady state. (i) Fragmented network, ($k_{u1}/k_{f}=30$) (ii) Network near percolation transition, with kinetic parameters appropriate to mammalian mitochondrial networks [57] ($k_{u1}/k_{f}=1000$). (iii) Hyperfused network ($k_{u1}/k_{f}=3000$). The ratio of tip-tip and tip-side fusion is set to $k_{u1}=3k_{u2}$, and particle diffusivity is $D_{p}=4800$ throughout.
@@
We seek to quantify the rate of material dispersion through dynamic spatial networks ranging from the social to the physical regime. These limiting architectures, as well as a continuum of intermediate network connectivities, are encompassed by a simulation framework [57], illustrated in Fig. 1, consisting of $N_{0}$ interacting spherocylindrical units of length $\ell_{0}$, which move diffusively through a bounded spherical domain of radius $R$ with unit diffusivity $D_{1}$. Upon close encounters, pairs of units may fuse into larger clusters with rate constants $k_{u1}$ (for tip-tip fusions) or $k_{u2}$ (for tip-side fusions). Fission breaks connections at the nodes, with fission rate proportional to the number of attached edges ($k_{f}$ at degree-2 nodes and $1.5k_{f}$ at degree-3 nodes), in keeping with prior models [13]. At steady state, the system exhibits a characteristic mean cluster size $\langle n\rangle$ set by the balance of fusion and fission, and an associated cluster diffusivity $D_{n}$. Details of this structural model and parameterization are described in Ref. [57]; here we focus on the spreading of material throughout the resulting dynamic networks.
@@
We select at random a single unit, which serves as a particle source with fixed constant concentration $h_{0}=1/\ell_{0}$. Such a fixed source could represent, for instance, an import site for nucleus-encoded proteins [58], a site of regulated translation at a mitochondrial RNA granule or nucleoid [59], or a location of calcium entry at a contact with the endoplasmic reticulum [60]. The concentration field is propagated between connected units on the network with effective diffusivity $D_{p}$, using a finite volume approach [61] discretized at the level of individual units. Concentrations decay with a constant rate $k_{d}$ (representing removal from the network). The decay rate serves to set a relevant timescale, with the resulting calculations closely related to the question of how much material spreads from the source over a certain time. We quantify dispersion by computing the total amount of material ($S$) in the network at steady-state, excluding the source unit. Example steady-state snapshots are shown in Fig. 1b and Supplemental Video 1.
@@
Given the complex internal structure of mitochondria and the variety of functionally relevant biomolecules within them, particle diffusivities can vary broadly. For example, proteins in the inner mitochondrial membrane can be embedded in the extensive folds of cristae [62]. The cristae also serve as diffusive barriers to hinder transport of solutes in the matrix [63] and intermembrane space [64]. Our simplified modeling approach coarse-grains these complications into a single effective diffusivity $D_{p}$ representing the motion of a particle along the mitochondrial tubule axis. Reported effective diffusivities vary over orders of magnitude, from $\sim 0.004\mu\text{m}^{2}/\text{s}$ for ATP synthase components  [65], to $\sim 20\mu\text{m}^{2}/\text{s}$ for mitochondrial matrix proteins [66], with diffusivity of ions presumed even higher [67]. We consider a range of diffusivities $D_{p}=0.4-40\mu\text{m}^{2}/\text{s}$ in our simulations. The diffusivity of individual mitochondria is estimated to be much slower, at $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$ [28, 68].
@@
The results below are reported in dimensionless units, relative to the length scale $2\ell_{0}$ and the timescale $1/k_{f}$. A reasonable estimate sets the mitochondrial unit length to $\ell_{0}=0.5\mu\text{m}$ and the fission timescale to $1/k_{f}\approx 2\text{ min}$ [57], consistent with experimental measurements of overall fission rate [69, 28]. The corresponding dimensionless diffusivities are $D_{p}=48-4800$ for the material, and $D_{1}=0.5$ for individual mitochondria.
@@
Figure 2: Material spreading on static networks of different dimension $d$. Two different sets of networks are considered: near-linear structures (blue) with $k_{u2}=0.01k_{u1}$ and highly-branched structures (green) with $k_{u2}=10k_{u1}$. The simulations are run to steady state, and the network structures are then frozen. (a) Calculation of the network dimension for a single instance of each network type (shown in inset). The number of nodes within a given graph distance is plotted against the graph distance on log-log axes, with the slope giving the dimension, $d$. Averaging over 21 snapshots from 3 independent simulations yields effective dimensionalities $d=1.2,d=1.9$, respectively. (b) Total material delivered from a source unit on the static networks is plotted as a function of the diffusive lengthscale, $\lambda_{p}=\sqrt{D_{p}/k_{d}}$. Solid curves show exact solution (details in SI Appendix), averaged over replicate snapshots. Dashed-circle curves show approximation with the linear motif (inset i), applicable at short $\lambda_{p}$. Dashed-square curves show continuum solution on a fractal domain (inset ii), matching well at long $\lambda_{p}$.
@@ Results
Spreading on the network is governed by the interplay of several key timescales: the decay time $\tau_{\text{d}}=1/k_{d}$, the cluster filling time $\tau_{\text{c}}$ for particles to diffusively explore a typical-sized connected component (cluster), and the waiting time between interactions $\tau_{\text{int}}$ (the time to encounter and fuse with a new cluster). The fission time $\tau_{f}=1/k_{f}$ sets the typical duration of transient interactions between distinct clusters. The material exhibits different dynamics depending on the comparative values of these four timescales. We begin by considering several limiting regimes where different timescales dominate.
@@
In the limit where $\tau_{d}$ is the shortest timescale, material spreads primarily within a single connected component. This regime is relevant for highly fused networks with large cluster sizes. Such networks tend to be relatively static in their topology, with any fissions that occur rapidly followed by re-fusion with the same neighboring unit [57]. Recent evidence of collocalizing mitochondrial fusion and fission machinery at ER-mitochondria contact sites also points towards increased likelihood of re-fusion following a fission event [70].
@@
For static networks, the steady-state distribution of material can be found by solving the diffusion equation on each one-dimensional edge, while matching boundary conditions at each node (details in SI Appendix). This method provides an exact solution for each individual network structure. However, to gain insight on the relevant scaling regimes, we turn to a mean-field continuum approach.
Depending on the relative rates of tip-tip versus tip-side fusion, network architectures can range from near-linear snake-like structures to highly branched compact morphologies, with the latter allowing for more rapid spreading of material through the network (Fig. 2). These structural features can be described by an effective fractal dimension $d$, which sets the scaling relationship between the number of units and the network ‘size’ in terms of graph distance. Specifically, we count the number of nodes within a given graph distance from a starting node, averaged over all possible starting nodes [71]. The power-law scaling exponent of this curve defines the dimension $d$ (Fig. 2i, inset) and therefore the cluster filling timescale, $\tau_{c}=(\left<n\right>^{1/d}\ell_{0})^{2}/D_{p}$. Although the simulated networks are embedded in a three-dimensional space, the limited junction degree gives rise to networks that are of lower dimension ($1\leq d\leq 2$). This feature has been previously observed in a variety of physical networks, from plant roots to ant tunnels, composed of tube-like objects connected at junctions [12].
@@
We first consider the limit where the diffusive lengthscale $\lambda_{p}=\sqrt{D_{p}/k_{d}}$ is smaller or comparable to the length of linear segments between junctions. In the case of a completely linear snake-like network, the steady-state material content $S_{1}$ is given by the solution of the one-dimensional (1D) diffusion equation, $S_{1}=\frac{n_{b}\lambda_{p}}{\ell_{0}}\tanh\left(\frac{\ell_{0}\left<n\right>/n_{b}}{\lambda_{p}}\right)$ with $n_{b}=2$. When there are additional branches adjacent to the source ($d>1$), more material is able to enter the network. We thus define $n_{b}$ as the average number of edges directly connected to the source unit, and approximate the entire network cluster as a set of $n_{b}$ linear spokes, each of length $\ell_{0}\left<n\right>/n_{b}$, connected to a central source (Fig. 2iviv). $S_{1}$ then gives an estimate of the total material spreading over a short diffusive lengthscale $\lambda_{p}$ (Fig. 2, dashed-circle curves).
@@
For the regime where material spreads over a more extensive network structure ($\lambda_{p}$ much greater than segment length), we consider diffusion through a $d$-dimensional medium. Here we approximate the spreading as spherically symmetric, with a $d$-sphere of radius $a_{n}$ (enclosing the source edge) maintained at fixed concentration, and a reflecting $d$-sphere of radius $R_{n}$ representing the outer boundary of the cluster. We set $R_{n}$ such that the average distance between two points in the $d$-dimensional domain is equal to the average graph distance between network nodes (details in SI Appendix). Assuming that the $\left<n\right$> network units are uniformly distributed within the continuum sphere, the radius $a_{n}$ is set proportionately to allow for 1 source unit within the inner sphere: $1/a_{n}^{d}=(\langle n\rangle-1)/(R_{n}^{d}-a_{n}^{d})$.
@@
We can write down the diffusion equation in $d$-space, $\left(\frac{1}{d}D_{p}\right)\frac{1}{r^{d-1}}\frac{\partial}{\partial r}\left(r^{d-1}\frac{\partial c(r)}{\partial r}\right)-k_{d}c(r)=0$, where the effective particle diffusivity is scaled by the dimension $d$. This scaling is applied because, when diffusing along an edge, the particle only moves in one dimension at each instance in time rather than simultaneously in all $d$ dimensions. The $d$-space diffusion equation has solutions in terms of modified Bessel functions $I_{\nu}(x),K_{\nu}(x)$ [72], where $\nu=1-d/2$ and $x=r\sqrt{d}/\lambda_{p}$. Taking a fixed concentration boundary at $x_{1}=a_{n}\sqrt{d}/\lambda_{p}$ and a reflecting boundary at $x_{2}=R_{n}\sqrt{d}/\lambda_{p}$ we find the total material delivered (details in SI Appendix):
[…]
@@
Fig. 2 shows the correspondence between diffusion through a fractal continuum (dashed-square curves) and the exact solution (solid curves) for networks with different connectivities. The linear motif solution $S_{1}$ is a good estimate for small diffusive lengths, while the continuum solution $S_{d}$ is a better approximation for large $\lambda_{p}$. When the diffusive lengthscale is large enough ($\lambda_{p}>R_{n}$), the entire connected component saturates. Overall, material delivery is boosted by increased network branching [21], which increases both the network dimension and the network density (lower $R_{n}$ for the same number of network units).

### v3

@@ Abstract
In eukaryotic cells, mitochondria form networks that range from highly fused interconnected structures to fragmented populations of individual organelles that undergo transient interactions. These structures can be described as temporal networks of physical units, whose dynamic topology is determined by fusion, fission, and motion of the mitochondria through intracellular space. The heterogeneity of the mitochondrial population is governed by diffusive transport and inter-unit exchange of proteins, lipids, ions, and RNA within these networks. We present a unifying framework for the dispersion of material within temporal networks of spatially embedded units that span across a broad connectivity range. Specifically, we consider filling of the networks with a locally produced but globally consumed material, demonstrating that the steady-state content is determined by the balance of timescales for spatial encounter between clusters, local fusion, fission, and diffusive transport within a cluster. As the connectivity increases, filling behavior transitions from three-dimensional spread through a ‘social network’ limited by cluster interactions to low-dimensional transport through a largely stationary ‘physical network’ limited by material diffusivity. We extract parameters for mitochondrial networks in three human cell lines, demonstrating that different cells can access both the social and the physical network regimes. These results provide a quantitative basis for predicting the homogenization of biomolecules through a mitochondrial population. Our framework unifies a variety of temporal network structures into an overarching theory for transport through populations of interacting and interconnected units.
@@
Diffusive transport through networks has been studied in a variety of contexts, including disease spread in epidemiology [1, 2, 3, 4], innovations in social networks [5], communication among insects [6, 7], diffusion of signals and nutrients in the brain extracellular space [8, 9], and oil recovery in porous rock [10]. These systems can be separated into two broad classes: one where the topology of network connections is stationary and limited to nearest neighbors, and another where network nodes are mobile and promiscuous, interacting with many different partners over time [4].
@@
Physical networks form an important category of topologically stationary, spatially-constrained network structures [11]. In such networks, edges represent objects subject to physical limitations (space-filling, steric repulsion, etc.) [12, 13] connecting degree-1 tips and degree-3 junctions, with higher degree nodes exceedingly rare [14, 15]. Examples include fungal mycelia [16], neuronal synaptic networks [17], and interconnected pores in rocks [18]. Transport behaviors on stationary networks can be described by the graph Laplacian, whose eigenvalues govern the spreading timescales [19, 20, 21]. In some percolation problems, physical networks become dynamic as edges are allowed to flicker between an active and inactive state [22, 23, 24]. However, the set of possible connections for each node remains tightly limited to its physically proximal neighbors.
@@
A distinct set of approaches considers signal transmission in ‘social’ networks of transiently-interacting units. In such networks, spreading is governed by distributions of contact duration and inter-contact times [6, 25], which may be constrained by spatial embedding [2, 1]. Models that use locally-interacting mobile agents moving through two-dimensional space have been shown to accurately represent topological properties for human friendship and sexual interaction networks [26]. Both the contact rate and the probability of transmission during each contact may depend on the particle mobility, resulting in multiple distinct scaling regimes as the mobility is increased [3]. Notably, many previously studied social network systems focus on the spreading of a non-diluting signal (infection, information, etc.) [6, 2, 5, 3, 4], in contrast to dispersion of mass-conserving physical material [27, 24, 28].
Despite the distinct modeling approaches employed, physical and social networks lie on a continuum of temporal network structures [25] with varying timescales of topological rearrangement. Understanding how the physical behavior of moving agents translates into temporal network properties and the transmission of signals upon such networks remains an open question in the field. Recent results have highlighted the importance of relative timescales for network rearrangement versus local transmission, with infection spread transitioning from a static locally-connected regime to a homogeneously mixed mean-field regime as the particle mobility increases [4].
@@
The intracellular environment exemplifies the full spectrum of temporal network regimes. For example, the endoplasmic reticulum forms a well-connected physical lattice of membrane-bound tubules that enables diffusive transport of proteins [28] and ions [29] across the entire cell. At the opposite extreme, the population of endocytic vesicles constitutes a social network of discrete transiently-interacting compartments [30, 31].
Mitochondria form another intracellular network with striking structural variability. Mitochondrial network architectures vary across cell types [32, 33, 34, 35], and transition between fragmented and hyperfused states in response to disease [36, 37, 38, 39, 40, 41], metabolic conditions [42, 43, 44], calcium signaling and apoptosis [45, 46], and cell cycle progression [47, 48]. Recent studies have highlighted the importance of mitochondrial dynamics in allowing the cell to tune between conflicting functional demands [49, 22]. On the one hand, maintaining a broad distribution of mitochondria throughout the cell helps limit ATP gradients [50, 51], facilitates the formation of contact sites with many other organelles [52], and allows mitochondria to serve as local calcium reservoirs and signal transduction hubs [53, 54]. On the other hand, fusion into more compact interconnected structures may contribute to reduction of mitochondrial heterogeneity [55, 56], genetic complementation [57, 58, 59], quality control of the mitochondrial population [60, 61, 62, 63], dilution of harmful reactive oxygen species (ROS) [64], and power cabling for enhanced ATP production [22, 65]. Rapid transport and transient fusion between small mitochondrial units, as observed in plant cells [35, 59, 49], enable broad dispersion of mitochondria throughout the cellular space while also allowing some mixing of mitochondrial contents within the population. The homogenization of mitochondrial material relies both on diffusive spreading within connected networks [66, 24] and on dynamic encounters and exchange events between disjoint clusters [49]. Establishing the functional consequences of different mitochondrial morphologies thus requires an overarching picture of the relative contributions from these distinct dynamic processes.
@@
To bridge the gap between mitochondrial structure and material transport, we present a quantitative framework for diffusion on spatially embedded dynamic networks. Our central finding is a connectivity-driven transition in transport: from low-dimensional tunneling through well-connected networks to three-dimensional dispersion across a population of transiently interacting units. We present a unifying expression that approximates the rate of network filling from a source across multiple distinct regimes, tuned by the interplay of competing timescales: the cluster filling rate, encounter rate, fusion rate, and the decay rate of the spreading material. The model is applied to mitochondrial networks in three human cell types, demonstrating the structural diversity and different regimes of relevance for these networks.
@@
Figure 1: Dynamic network simulation framework. (a) Schematic of network formation by interacting units which diffuse through 3D space with diffusivity $D_{1}$. Nearby units can undergo tip-tip and tip-side fusion with rates $k_{u1},k_{u2}$ while connected nodes undergo fission at rate $k_{f},\frac{3}{2}k_{f}$, respectively. Material spreads along connected units with diffusivity $D_{p}$ and decays over time with rate $k_{d}$. Red indicates the concentration of material in a given unit. (b) Snapshots of network filling from a single source unit, at steady state. (i) Fragmented network, ($k_{u1}/k_{f}=30$) (ii) Network near percolation transition, with kinetic parameters appropriate to mammalian mitochondrial networks [67] ($k_{u1}/k_{f}=1000$). (iii) Hyperfused network ($k_{u1}/k_{f}=3000$). The ratio of tip-tip and tip-side fusion is set to $k_{u1}=3k_{u2}$, the dimensionless particle diffusivity is $D_{p}=4800$, and the dimensionless material decay rate is $k_{d}=0.32$ throughout.
@@
We seek to quantify the rate of material dispersion through dynamic spatial networks ranging from the social to the physical regime. These limiting architectures, as well as a continuum of intermediate network connectivities, are encompassed by a simulation framework [67], illustrated in Fig. 1, consisting of $N_{0}$ interacting spherocylindrical units of length $\ell_{0}$. Each node (unit endpoint) is subject to random, uncorrelated Brownian forces, resulting in an effective translational diffusivity $D_{1}$ for individual units. The nodes are also subject to elastic bending, stretching, and steric repulsion forces [67], analogous to standard polymer mechanics models. Upon close encounters, pairs of units may fuse into larger clusters with rate constants $k_{u1}$ (for tip-tip fusions) or $k_{u2}$ (for tip-side fusions). Fission breaks connections at the nodes, with fission rate proportional to the number of attached edges ($k_{f}$ at degree-2 nodes and $1.5k_{f}$ at degree-3 nodes), in keeping with prior models [15]. The unit length $\ell_{0}$ thus represents the minimal cluster size that can be formed by fission and the minimal separation between junctions. At steady state, the system exhibits a characteristic mean cluster size $\langle n\rangle$ set by the balance of fusion and fission, and an associated translational cluster diffusivity $D_{n}$.
We note that this structural model differs from classic diffusion-limited aggregation or cluster-cluster aggregation models [68, 69] in that the degree at each node is limited, there are mechanical moduli that constrain bending at linear nodes and junctions, and the fusion rate is dependent on angle and node degree. The spherocylindrical units enable explicit incorporation of these effects to represent the physical features of mitochondrial networks, including separately tuning the length of snake-like segments between junctions and the cluster size of the networks. Details of the structural model and the selection of appropriate parameters are described in Ref. [67], and summarized in the SI Appendix. Here we focus on the spreading of material throughout the resulting dynamic networks.
@@
The network structures are allowed to run to steady state prior to the start of the material spreading simulation. We then select at random a single unit, which serves as a particle source with fixed constant concentration $h_{0}=1/\ell_{0}$. Such a fixed source could represent, for instance, an import site for nucleus-encoded proteins [70], a site of regulated translation at a mitochondrial RNA granule or nucleoid [71], or a location of calcium entry at a contact with the endoplasmic reticulum [72]. The concentration field is propagated between connected units on the network with effective diffusivity $D_{p}$, using a finite volume approach [73] discretized at the level of individual units. Concentrations decay with a constant rate $k_{d}$ (representing removal from the network). The decay rate serves to set a relevant timescale, with the resulting calculations closely related to the question of how much material spreads from the source over a certain time. We quantify dispersion by computing the total amount of material ($S$) in the network at steady-state, excluding the source unit. Example steady-state snapshots are shown in Fig. 1b and Supplemental Video 1.
@@
Given the complex internal structure of mitochondria and the variety of functionally relevant biomolecules within them, particle diffusivities can vary broadly. For example, proteins in the inner mitochondrial membrane can be embedded in the extensive folds of cristae, resulting in much slower diffusivity than the corresponding outer membrane proteins [74, 75]. The cristae also serve as diffusive barriers to hinder transport of solutes in the matrix [76] and intermembrane space [77]. Our simplified modeling approach coarse-grains these complications into a single effective diffusivity $D_{p}$ representing the motion of a particle along the mitochondrial tubule axis. Reported effective diffusivities vary over orders of magnitude, from $\sim 0.004\mu\text{m}^{2}/\text{s}$ for ATP synthase components  [78], to $\sim 20\mu\text{m}^{2}/\text{s}$ for mitochondrial matrix proteins [79], with diffusivity of ions presumed even higher [80]. We consider a range of diffusivities $D_{p}=0.4-40\mu\text{m}^{2}/\text{s}$ in our simulations. The averaged diffusivity of individual mitochondria, which includes occasional active transport along microtubules, is estimated to be much slower, at $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$ [32, 81].
@@
The results below are reported in dimensionless units, relative to the length scale $2\ell_{0}$ and the timescale $1/k_{f}$. A reasonable estimate sets the mitochondrial unit length to $\ell_{0}=0.5\mu\text{m}$ and the fission timescale to $1/k_{f}\approx 2\text{ min}$ [67], consistent with experimental measurements of fragmented mitochondrial volume [82, 83, 84] and overall fission rate [85, 32]. The corresponding dimensionless diffusivities are $D_{p}=48-4800$ for the material, and $D_{1}=0.5$ for individual mitochondria.
@@
Figure 2: Filling of static networks depends on network dimensionality. Two different sets of networks are considered: near-linear structures (blue) with $k_{u2}=0.01k_{u1}$ and highly-branched structures (green) with $k_{u2}=10k_{u1}$. The simulations are run to steady state, and the network structures are then frozen. (a) Steady-state material content supplied by a source unit on the static networks is plotted as a function of the diffusive lengthscale, $\lambda_{p}=\sqrt{D_{p}/k_{d}}$. Solid curves show exact solution (details in SI Appendix), averaged over replicate snapshots. Dashed-circle curves show approximation with the linear motif (inset i), applicable at short $\lambda_{p}$. Dashed-square curves show continuum solution on a fractal domain (inset ii), applicable at long $\lambda_{p}$. (b) Calculation of the graph dimension for a single instance of each network type (shown in inset). The number of nodes within a given graph distance is plotted against the graph distance on log-log axes, with the slope giving the intrinsic dimension $d$. Averaging over 21 snapshots from 3 independent simulations yields effective dimensionalities $d=1.2,d=1.9$, respectively.
@@ Results
Spreading on the network is governed by the interplay of several key timescales: the decay time $\tau_{\text{d}}=1/k_{d}$, the cluster filling time $\tau_{\text{c}}$ for particles to diffusively explore a typical-sized connected component (cluster), the waiting time between interactions $\tau_{\text{int}}$ (the time to encounter and fuse with a new cluster), and the fission time $\tau_{f}=1/k_{f}$ which sets the typical duration of transient interactions between clusters. The material exhibits different dynamics depending on the comparative values of these four timescales. We begin by considering limiting regimes where different timescales dominate.
@@
In the limit where $\tau_{d}$ is the shortest timescale, material spreads primarily within a single connected component. This regime is relevant for highly fused networks with large cluster sizes. Such networks tend to be relatively static in their topology, with any fissions that occur rapidly followed by re-fusion with the same neighboring unit [67]. Recent evidence of colocalizing mitochondrial fusion and fission machinery at ER-mitochondria contact sites also points towards increased likelihood of re-fusion following a fission event [86].
@@
For static networks, the steady-state distribution of material can be found by solving the diffusion equation on each one-dimensional edge, while matching boundary conditions at each node (details in SI Appendix). This method yields an exact solution for each individual network (Fig. 2a). However, simplified approximations of the network structure provide further insight on the relevant scaling regimes.
@@
Depending on the relative rates of tip-tip versus tip-side fusion, network architectures can range from near-linear snake-like structures to highly branched compact morphologies, with the latter allowing for more rapid spreading of material through the network (Fig. 2). We first consider the snake-like limit, relevant when the diffusive length scale $\lambda_{p}=\sqrt{D_{p}/k_{d}}$ is smaller or comparable to the length of linear segments between junctions. In the case of a completely linear network, the steady-state material content $S_{1}$ is given by the solution of the one-dimensional (1D) diffusion equation, $S_{1}=\frac{n_{b}\lambda_{p}}{\ell_{0}}\tanh\left(\frac{\ell_{0}\left<n\right>/n_{b}}{\lambda_{p}}\right)$ with $n_{b}=2$. Moving beyond the linear limit, additional branches adjacent to the source enable more material to enter the network. We thus define $n_{b}$ as the average number of edges directly connected to the source unit, and approximate the entire network cluster as a set of $n_{b}$ linear spokes, each of length $\ell_{0}\left<n\right>/n_{b}$, connected to a central source (Fig. 2ii). $S_{1}$ then gives an estimate of the total material spreading over a short diffusive length-scale $\lambda_{p}$ (Fig. 2a).
For the regime where material spreads over a more extensive network structure ($\lambda_{p}$ much greater than segment length), we turn to a fractal continuum approach. Complex and porous media are often described as self-similar fractal structures characterized by polynomial scaling laws that relate different measures of distance and system size [87, 88, 89]. For our simulated network structures, we count the number of nodes within a given graph distance from a starting node, averaged over all possible starting nodes [90]. The power-law scaling exponent of this curve (Fig. 2b) defines the intrinsic fractal dimension $d$ of the network. Sometimes called $d_{\ell}$, the ‘chemical’ or ‘graph’ dimension [91, 92], this value describes the topology of the network and is distinct from the extrinsic dimension describing how the network is laid out in space [87, 93].
Although the simulated networks are embedded in three-dimensional space, their graph dimension is substantially lower ($1\leq d\leq 2$). This feature has been previously observed in a variety of physical networks, from plant roots to ant tunnels, composed of tube-like objects connected at junctions [14]. Three-dimensional percolation networks, often used to describe transport through porous media [94], have a graph dimension of $d\approx 1.8$ [87, 93]. Structures arising from irreversible cluster-cluster aggregation have $d\approx 1.4$ [69]. These low dimensionalities imply that the tortuosity of paths [95, 96] (ratio of path length to Euclidean distance) increases with system size, leading to effectively subdiffusive motion through space for particles embedded within the network [93]. For a network of $n$ units, with intrinsic graph dimension $d$, the cluster filling timescale can be estimated as: $\tau_{c}=(\left<n\right>^{1/d}\ell_{0})^{2}/D_{p}$.
@@
To compute the steady-state material content, we approximate the network as a hyperspherical domain of dimension $d$, with a sphere of radius $a_{n}$ (enclosing the source edge) maintained at fixed concentration, and a reflecting sphere of radius $R_{n}$ representing the outer boundary of the cluster (Fig. 2iiii). We set $R_{n}$ such that the average distance between two points in the $d$-dimensional domain is equal to the average graph distance between network nodes (details in SI Appendix). Assuming that the $\left<n\right$> network units are uniformly distributed within the continuum sphere, the radius $a_{n}$ is set proportionately to allow for 1 source unit within the inner sphere: $1/a_{n}^{d}=(\langle n\rangle-1)/(R_{n}^{d}-a_{n}^{d})$.
@@
Assuming the system is spherically symmetric, we can write down the steady-state diffusion equation in $d$-dimensional space [97] in terms of the radial coordinate $r$, as $\left(\frac{1}{d}D_{p}\right)\frac{1}{r^{d-1}}\frac{\partial}{\partial r}\left(r^{d-1}\frac{\partial c(r)}{\partial r}\right)-k_{d}c(r)=0$. The effective particle diffusivity is scaled by the dimension $d$ because, when diffusing along an edge, the particle only moves in one dimension at each instance in time rather than simultaneously in all $d$ dimensions. The $d$-space diffusion equation has solutions in terms of modified Bessel functions $I_{\nu}(x),K_{\nu}(x)$ [97], where $\nu=1-d/2$ and $x=r\sqrt{d}/\lambda_{p}$. Taking a fixed concentration boundary at $x_{1}=a_{n}\sqrt{d}/\lambda_{p}$ and a reflecting boundary at $x_{2}=R_{n}\sqrt{d}/\lambda_{p}$ we find the steady-state material content (details in SI Appendix):
$gathered S_{d}=\frac{d}{x_{1}}\frac{\left[-I_{\nu-1}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu-1}(x_{1})\right]}{\left[I_{\nu}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu}(x_{1})\right]}\\$ (1)
@@
Fig. 2a shows the correspondence between diffusion through a fractal continuum and the exact solution for networks with different connectivities. The linear motif solution $S_{1}$ is a good estimate for small diffusive lengths, while the continuum solution $S_{d}$ is a better approximation for large $\lambda_{p}$. This transition is analogous to a ‘representative elementary volume’ used to describe the minimal size of porous media above which a self-similar fractal approximation can apply [98, 96]. When the diffusive length-scale is large enough ($\lambda_{p}>R_{n}$), the entire connected component saturates. Overall, material delivery is boosted by increased network branching [24], which increases both the network dimension and the network density (lower $R_{n}$ for the same number of network units).